A parent can choose from 7 types of protein, 5 vegeatables, and 3 desserts. if the parent serves a meal of 1 protein, 1 vegetable, and 1 dessert, how many different meals can be served?
Answer by jim_thompson5910(35256)

There are 7*5*3 = 35*3 = 105 different meals.